Deere & Company (DE) has a current stock price of $83.74 and is expected to pay a $2.04 dividend in one year. DE has a beta of 1.32, the expected market return is 7.5% and the risk-free rate is 3.0%. What is DE’s expected price in one year?

A.

$94.54

B.

$91.23

C.

$89.19

D.

$92.50
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

=3+1.32*(7.5-3)

=8.94%

Required return=(D1/Current price)+Growth rate

0.0894=(2.04/83.74)+Growth rate

Growth rate=0.0894-(2.04/83.74)

=0.0650388823

Hence expected price=Current price*(1+Growth rate)

=83.74*(1+0.0650388823)

=$89.19(Approx).